1. How long would it take Jesse with an acceleration of -2.50 m/s2 to bring his bicycle with an initial velocity of 13.5 m/s to a complete stop? 
1
Expert's answer
2022-03-16T18:28:21-0400

The required time

"\Delta t=\frac{\Delta v}{a}=\frac{0-13.5}{-2.50}=5.40\:\rm s"